# CAMPBELL & REMIRO-AZÓCAR

### v1

@@ 5 Applied example
We now demonstrate the application of some of the methods outlined in Section 3 to synthetic lung cancer clinical trial data. The data were obtained from the “MAIC” R package, implemented by researchers from the pharmaceutical industry.95 Our objective is to compare the objective response, a binary outcome $Y$, under two treatments: the active “intervention” ($T=1$) and the external “control” ($T=0$). The data consist of IPD from a SAT ($S=1$) with $n_{1}=500$ subjects under the “intervention”, and AD from an external historical SAT ($S=0$) of $n_{0}=300$, which makes up the “control”. The unavailability of IPD for the external control allows us to illustrate the methodological extensions described in Section 3.7. R code to reproduce the applied example is provided in the Supplementary Material.
@@
Covariate Intervention SAT External control Weighted intervention SAT
[…]
Age in years 59.85; 9.01 50.06; 3.24 50.06; 3.24
(mean; standard deviation)
Sex 0.38 0.49 0.49
(proportion male)
ECOG 0.41 0.35 0.35
(proportion ECOG performance status of 1)
Smoking status 0.32 0.19 0.19
(proportion of smokers)
@@
Table 5: Summary statistics of the four baseline covariates identified as imbalanced prognostic factors. Note that the standard deviation of age in the weighted intervention SAT is $\sum_{i=1}^{n_{1}}\left(\omega_{i}(X_{1,i}-\sum_{i=1}^{n_{1}}\omega_{i}X_{1,i})^{2}\right)$, where $X_{1,i}$ and $\omega_{i}$ are the age and the MAIC weight, respectively, for subject $i=1,\dots,n_{1}$ in the intervention SAT.
@@
We consider the naïve estimator first, which does not perform covariate adjustment (Equation 35). In the intervention SAT, 390 of $n_{1}=500$ subjects attained objective response, which equates to a 78% response rate, $\hat{\mu}_{1}^{1}=0.78$. This implies a log-odds of response of $g(\hat{\mu}_{1}^{1})=1.266$, where $g(\cdot)=\operatorname{logit}(\cdot)$. In the external control, 120 of $n_{0}=300$ subjects attained objective response, which equates to a 40% response rate, $\hat{\mu}_{0}^{0}=0.4$. This implies a log-odds of response of $g(\hat{\mu}_{0}^{0})=-0.405$. A naïve estimate is obtained by simple subtraction:
$\widehat{\textrm{ATC}}_{naive}=1.266-(-0.405)=1.671.$
Using the Delta method, we obtain:
$\textrm{SE}\left(g(\hat{\mu}_{1}^{1})\right)$ $=\sqrt{{1}/{500(0.78\times(1-0.78))}}$
$=0.108,$
$\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)$ $=\sqrt{{1}/{300(0.40\times(1-0.40))}}$
$=0.118.$
Then, following Equation 33, we have:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{naive}\right)$ $=\sqrt{(\textrm{SE}\left(g\left(\hat{\mu}_{1}^{1})\right)\right)^{2}+(\textrm{SE}\left(g\left(\hat{\mu}_{0}^{0})\right)\right)^{2}}$
$=\sqrt{0.108^{2}+0.118^{2}}$
$=0.160.$
Based on the computed standard errors, Wald-type 95% confidence intervals are given by:
$(\widehat{\textrm{ATC}}_{naive}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{naive}),\widehat{\textrm{ATC}}_{naive}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{naive}))=(1.358,1.984)$.
@@
We now consider the MAIC estimator. The positivity assumption is assessed using a method proposed by Glimm and Yau (2022), which verifies whether covariate AD from the external control lie within the convex hull of the SAT covariate space, and whether the MAIC numerical optimization algorithm can converge.60 This method is implemented using the “maicLP” R function in the “maicChecks” R package,96 which confirms that a feasible weighting solution to the MAIC convex optimization problem exists, i.e., that there is a set of positive weights that can enforce covariate balance between the intervention SAT and the external control, and that the MAIC numerical optimization algorithm can converge.
@@
where $\mathbf{c(X)}$ is a 5-by-500 matrix, with the rows representing the age, sex, ECOG performance status, smoking status, and age-squared for subjects in the intervention SAT. Using the BFGS convex optimization algorithm to minimize the objective function in Equation 13, we obtain $\mathbf{\hat{\gamma}}=(3.542,0.589,-0.698,-0.048,-0.036)$, and weights are calculated subject to the constraint that they sum to one. Figure 1 shows a histogram illustrating the empirical distribution of the resulting MAIC weights. The effective sample size (ESS) – that is, the number of independent non-weighted observations that would be required to give an estimate with approximately the same precision as the weighted sample estimate – of the intervention SAT after weighting is:
$\textrm{ESS}=\frac{\left(\sum_{i=1}^{n_{1}}\omega_{i}\right)^{2}}{\sum_{i=1}^{n_{1}}\left(\omega_{i}^{2}\right)}=\frac{1}{0.0063665}=157.07.$
Figure 1: Histogram of the EB (MAIC) weights; EB denotes entropy balancing.
Using the MAIC weights, the ATC is estimated as:
$\widehat{\textrm{ATC}}_{MAIC}=$ $g\underbrace{\left(\sum_{i=1}^{n_{1}}\hat{\omega}_{i}Y_{i}\right)}_{\hat{\mu}_{0}^{1}}-\hat{\mu}_{0}^{0}$
$=0.926-(-0.405)$
$=1.331.$
We previously calculated $\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)=0.118$. Using the non-parametric bootstrap with $B=10000$ resamples, $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.177$. As per Equation 33:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{MAIC}\right)$ $=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}}$
$=\sqrt{0.177^{2}+0.118^{2}}$
$=0.212,$
and Wald-type 95% confidence intervals are given by
$(\widehat{\textrm{ATC}}_{MAIC}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{MAIC}),\widehat{\textrm{ATC}}_{MAIC}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{MAIC}))=(0.915,1.748)$.
To perform G-computation with unavailable IPD for the external control, we first simulate $M=10000$ individual-level covariate profiles from an assumed joint covariate distribution for the external control, as per the approach outlined by Remiro-Azócar et al (2022)46 or the “infinite population” STC method described by Zhang et al (2024).97 We proceed by assuming that the pairwise correlation structure of the four covariates in the external control is equal to that observed in the intervention SAT:
@@
with the rows/columns in the order: age, sex, ECOG performance status and smoking status. For age, we assumed a Normal(50.06, 3.24) marginal distribution; for sex, ECOG performance status and smoking status, we assumed Bernoulli(0.49), Bernoulli(0.35) and Bernoulli(0.19) marginal distributions, respectively, based on the summary statistics of the external control in Table 5. Individual-level covariates were ultimately simulated from a Gaussian copula characterized by the aforementioned marginal distributions and pairwise correlation structure, using the “add_integration” function from the multinma R package.98
@@
for $i$ in $n_{1}$, where $\mathbf{\beta}=\left(\beta_{0},\beta_{1},\beta_{2},\beta_{3},\beta_{4},\beta_{5}\right)^{\top}$ are regression coefficients. The model was fitted to the intervention SAT using maximum-likelihood estimation with Fisher scoring, and we obtained regression coefficient estimates of $\hat{\beta}_{0}=5.72$, $\hat{\beta}_{1}=-0.20$, $\hat{\beta}_{2}=0.12$, $\hat{\beta}_{3}=0.13$, $\hat{\beta}_{4}=0.01$, and $\hat{\beta}_{5}=0.00$.
Then, following the G-computation procedure outlined in Section 3.4, the ATC is estimated as:
$\widehat{\textrm{ATC}}_{Gcomp}$ $=g\underbrace{\left(\frac{1}{M}\sum_{i=n_{1}+1}^{n_{1}+M}\hat{Y}_{i}^{1}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right)$
$=g{\left(\frac{7149.896}{10000}\right)}-(-0.405)$
$=1.325,$
where $\hat{Y}_{i}^{1}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$ for $i=n_{1}$,…,$n_{1}+M$, are predictions of the fitted outcome model – indexed $m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})$ with regression coefficient point estimates $\hat{\mathbf{\beta}}$ – for each of the $M$ simulated individual-level covariate profiles.
Using the non-parametric bootstrap with $B=10000$ resamples, $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.164$. As per Equation 33:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{Gcomp}\right)$ $=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}}$
$=\sqrt{0.164^{2}+0.118^{2}}$
$=0.202,$
and Wald-type 95% confidence intervals are given by
$(\widehat{\textrm{ATC}}_{Gcomp}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{Gcomp}),\widehat{\textrm{ATC}}_{Gcomp}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{Gcomp}))=(0.929,1.722)$.
Finally, our doubly robust augmented MAIC estimator proposed in Section 3.4 produces the estimate:
@@
$=g\underbrace{\left(0.001+\frac{7149.896}{10000}\right)}_{\hat{\mu}_{0}^{1}}-g\left(\hat{\mu}_{0}^{0}\right)$
$=0.926-(-0.405)$
$=1.332,$
where $\hat{Y}_{i}^{1}=q^{-1}\left(m({\mathbf{X}_{i}};\hat{\mathbf{\beta}})\right)$ are predictions of the fitted outcome model for each of the subjects in the intervention SAT ($i=1,...,n_{1}$) and each of the $M$ simulated individual-level covariate profiles ($i=n+1,\dots,n_{1}+M$).
Using the non-parametric bootstrap with $B=10000$ resamples, $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.179$. As per Equation 33:
$\textrm{SE}\left(\widehat{\textrm{ATC}}_{DR}\right)$ $=\sqrt{\big{(}\textrm{SE}\left(g(\hat{\mu}^{1}_{0})\right)\big{)}^{2}+\big{(}\textrm{SE}\left(g(\hat{\mu}^{0}_{0})\right)\big{)}^{2}}$
$=\sqrt{0.179^{2}+0.118^{2}}$
$=0.214,$
and Wald-type 95% confidence intervals are given by
$(\widehat{\textrm{ATC}}_{DR}-1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{DR}),\widehat{\textrm{ATC}}_{DR}+1.96\times\textrm{SE}(\widehat{\textrm{ATC}}_{DR}))=(0.912,1.751)$.
Figure 2 shows the point estimates obtained using the four different estimators alongside their 95% confidence intervals. When comparing the covariate-adjusted approaches to the naïve approach, we observe that covariate adjustment shifts the point estimate towards the null considerably.
Figure 2: Point estimates with 95% confidence intervals of the ATC (marginal log-odds ratio of objective response) for the different estimators in the applied example. DR denotes doubly robust and EB denotes entropy balancing.
When comparing MAIC, G-computation and our proposed doubly robust (DR) augmented MAIC estimator, results across the three methods seem consistent. Despite the shift towards the null, results suggest that the intervention improves objective response versus the control, statistically significant at the 5% level. In this case, the DR point estimate is not meaningfully different than the MAIC or G-computation point estimates, and the DR approach offers slightly increased standard errors and wider confidence intervals than G-computation. Nevertheless, this loss of precision seems a relatively minor price to pay, compensated for by greater reassurance in our results due to increased protection against misspecification of the outcome model.
@@ 6 Discussion
The objective of this article was to clarify existing approaches for doubly robust estimation in the context of externally controlled SATs, and to propose a novel estimator that augments approaches based on MAIC or entropy balancing via an outcome model. We described and illustrated an extension of this estimator specifically tailored to unanchored ITCs, for the setting with unavailable external control IPD, which is commonly encountered in practice. In a simulation study and applied example, we evaluated the performance and demonstrated the use of different doubly robust augmented estimators, highlighting their merits with respect to the more popular non-augmented singly robust estimators.
@@
Our findings reinforce the understanding that “balancing” approaches to weighting, such as MAIC or entropy balancing, can enhance performance relative to standard “modeling” approaches, but lack the doubly robust property for non-linear outcome models. Conversely, the augmented MAIC or entropy balancing-based estimator demonstrates double robustness with a logistic outcome model, and exhibits higher precision than non-augmented weighting estimators when the outcome model is correctly specified. Moreover, it generally achieves near-identical precision to G-computation, which offers the lowest variance under correct specification of the outcome model, but may exhibit notable bias where the outcome model is misspecified.
@@
A potential concern about augmented estimators has been possible bias and variance amplification where both the propensity score and the outcome model are incorrectly specified. Nevertheless, such amplification under dual model misspecification was not observed in our simulation study, and one can argue that risks are mitigated by employing “balancing” instead of “modeling” weights (for reasons outlined in the introduction to Section 3.3), but further theoretical work and simulation studies are required to fully support this claim. The results of our simulation study motivate the routine application of doubly robust augmented estimators, particularly those based on MAIC or entropy balancing weights, in practical settings. It is unfortunate to see that virtually all applications of MAIC do not consider augmentation via an outcome model, and that most practical uses of augmented estimators apply “modeling” instead of “balancing” weights.99
@@
While the “weighted G-computation” estimator described in Section 3.5 exhibited double robustness for the ATC in our simulation study and almost identical performance to our proposed augmented MAIC estimator, these results are only expected to hold where the outcome model is a GLM with canonical link function, and not where the link function is non-canonical or for outcome models in the time-to-event setting.51 Future simulation studies should consider non-binary outcomes, including survival outcomes with censoring, and other summary effect measures beyond the log-odds ratio.
@@
Our simulation study and applied example considered scenarios with relatively low sample sizes in the SAT and the external control, corresponding to typical settings in rare disease and late-stage hematological or solid tumor oncology, where the number of subjects enrolled in SATs can be one- or two-hundred, but may also consist of several hundreds. The number of external controls can be equally small; as such, our findings are potentially applicable where the target estimand is the ATT instead the ATC, which requires weighting the external control as opposed to the SAT. Somewhat worryingly, all augmented and non-augmented weighting estimators displayed some small-sample bias in our simulation study under a total sample size of $n=200$, even if modeling assumptions were correct.
@@
The level of (deterministic) overlap between the SAT and external control covariate distributions in our simulation study was relatively strong. The performance of augmented and non-augmented weighting estimators with respect to G-computation will likely worsen as overlap decreases, particularly in conjunction with small sample sizes. Nevertheless, we conjecture that the performance of balancing-based approaches will suffer less than that of their corresponding modeling-based counterparts, due to generating more stable and less extreme weights. This is unless a complete lack of overlap results in the absence of a solution to the convex optimization balancing problem, in which case extrapolating via G-computation would be the only option.
@@
Finally, it is important to emphasize that all the covariate-adjusted estimators we have considered, including the doubly robust augmented approaches, require the important assumption of no unmeasured prognostic factors, and can incur bias if these are omitted or missing. In practice, important prognostic factors may be unknown or unavailable in at least one of the SAT or external control data sources. An important area of future research is the development of sensitivity analysis or quantitative bias analysis methods to help explore the sensitivity of results to unmeasured prognostic factors, in the specific context of externally controlled SATs and unanchored ITCs.100

### v2

@@ 5 Applied example
We now demonstrate the application of some of the methods outlined in Section 3 to synthetic lung cancer clinical trial data. The data were obtained from the “MAIC” R package, implemented by researchers from the pharmaceutical industry.iain_t_bennett_roche_2022_6624152 Our objective is to compare the objective response, a binary outcome $Y$, under two treatments: the active “intervention” ($T=1$) and the external “control” ($T=0$). The data consist of IPD from a SAT ($S=1$) with $n_{1}=500$ subjects under the “intervention”, and AD from an external historical SAT ($S=0$) of $n_{0}=300$, which makes up the “control”. The unavailability of IPD for the external control allows us to illustrate the methodological extensions described in Section 3.8. R code to reproduce the applied example is provided in the Supplementary Material.
@@
Covariate Intervention SAT External control Normalized IOW MAIC
weighted SAT weighted SAT
($n_{1}=500$) ($n_{0}=300$) ($\textrm{ESS}=153.42$) ($\textrm{ESS}=157.07$)
Age in years (mean; SD) 59.85; 9.01 50.06; 3.24 49.53; 3.18 50.06; 3.24
Sex (proportion male) 0.38 0.49 0.45 0.49
ECOG (proportion status 1) 0.41 0.35 0.30 0.35
Smoking (proportion smokers) 0.32 0.19 0.20 0.19
@@
Table 5: Summary statistics of the four baseline covariates identified as imbalanced prognostic factors, before and after weighting using MAIC (entropy balancing) and normalized inverse odds weighting (IOW). The standard deviation of age in the weighted columns is $\sqrt{\sum_{i=1}^{n_{1}}v_{i}(X_{1,i}-\sum_{i=1}^{n_{1}}v_{i}X_{1,i})^{2}}$, where $X_{1,i}$ and $v_{i}$ are the age and the weight, respectively, for subject $i=1,\dots,n_{1}$ in the intervention SAT.
@@
We consider the naïve estimator first, which does not perform covariate adjustment (Equation 27). In the intervention SAT, 390 of $n_{1}=500$ subjects attained objective response, which equates to a 78% response rate, $\hat{\mu}_{1}^{1}=0.78$. This implies a log-odds of response of $g(\hat{\mu}_{1}^{1})=1.266$, where $g(\cdot)=\operatorname{logit}(\cdot)$. In the external control, 120 of $n_{0}=300$ subjects attained objective response, which equates to a 40% response rate, $\hat{\mu}_{0}^{0}=0.40$. This implies a log-odds of response of $g(\hat{\mu}_{0}^{0})=-0.405$. A naïve estimate is obtained by simple subtraction: $\widehat{\textrm{ATC}}_{naive}=1.266-(-0.405)=1.671$ and using the Delta method, we obtain: $\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)=0.118$ and $\textrm{SE}\left(g(\hat{\mu}_{1}^{1})\right)=0.108$, and a Wald-type 95% CI of $(1.358,1.984)$.
Next, consider the normalized IOW estimator. A logistic regression model for the probability of SAT participation, conditional on age, sex, ECOG performance status, smoking status, and age-squared is fitted to the concatenated individual-level data comprising the intervention SAT and $M=10000$ simulated covariate profiles for the external control. The estimated propensity scores are used to derive inverse odds weights, which are subsequently normalized to sum to one, yielding the Hajek-type estimator in Equation 4. Table 5 shows that the IOW-weighted covariate means are close to, but do not exactly match, those of the external control. The effective sample size of the IOW-weighted intervention SAT is 153.42 and standardized mean differences are substantially reduced after weighting (e.g., from $1.445$ to 0.079-0.079 for age), indicating that the weights achieve adequate, though not exact, balance. Using the normalized IOW weights, the ATC is estimated as: $\widehat{\textrm{ATC}}_{IOW}=1.333$. Having previously calculated $\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)=0.118$, we use the non-parametric bootstrap with $B=10000$ resamples to obtain $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.183$, and per Equation 26, we then obtain $\textrm{SE}\left(\widehat{\textrm{ATC}}_{IOW}\right)=0.216$ and a Wald-type 95% CI of $(0.911,1.756)$.
@@
We now consider the MAIC estimator. The positivity assumption is assessed using a method proposed by Glimm and Yau (2022), which verifies whether covariate AD from the external control lie within the convex hull of the SAT covariate space, and whether the MAIC numerical optimization algorithm can converge.glimm2022geometric This method is implemented using the “maicLP” R function in the “maicChecks” R package,maicchecks which confirms that a feasible weighting solution to the MAIC convex optimization problem exists, i.e., that there is a set of positive weights that can enforce covariate balance between the intervention SAT and the external control, and that the MAIC numerical optimization algorithm can converge.
@@
where $\mathbf{c(X)}$ is a 5-by-500 matrix, with the rows representing the age, sex, ECOG performance status, smoking status, and age-squared for subjects in the intervention SAT. Using the BFGS convex optimization algorithm to minimize the objective function in Equation 10, we obtain $\mathbf{\hat{\gamma}}=(3.542,0.589,-0.698,-0.048,-0.036)$, and weights are calculated subject to the constraint that they sum to one. Figure 1 shows a histogram illustrating the empirical distribution of the resulting MAIC weights compared to the normalized IOW weights. The effective sample size (ESS) – that is, the number of independent non-weighted observations that would be required to give an estimate with approximately the same precision as the weighted sample estimate – of the intervention SAT after weighting is 157.07. Using the MAIC weights, the ATC is estimated as: $\widehat{\textrm{ATC}}_{MAIC}=1.331$. Having previously calculated $\textrm{SE}\left(g(\hat{\mu}_{0}^{0})\right)=0.118$, we use the non-parametric bootstrap with $B=10000$ resamples to obtain $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.177$, and per Equation 26, we then obtain $\textrm{SE}\left(\widehat{\textrm{ATC}}_{MAIC}\right)=0.212$ and a Wald-type 95% CI of $(0.915,1.748)$.
Figure 1: Histogram of the normalized IOW weights (left) and MAIC (entropy balancing) weights (right).
To perform G-computation with unavailable IPD for the external control, we first simulate $M=10000$ individual-level covariate profiles from an assumed joint covariate distribution for the external control, as per the approach outlined by Remiro-Azócar et al (2022)remiro2022parametric or the “infinite population” STC method described by Zhang et al (2024).zhang2024four We proceed by assuming that the pairwise correlation structure of the four covariates in the external control is equal to that observed in the intervention SAT:
@@
with the rows/columns in the order: age, sex, ECOG performance status and smoking status. For age, we assumed a Normal(50.06, 3.24) marginal distribution; for sex, ECOG performance status and smoking status, we assumed Bernoulli(0.49), Bernoulli(0.35) and Bernoulli(0.19) marginal distributions, respectively, based on the summary statistics of the external control in Table 5. Individual-level covariates were ultimately simulated from a Gaussian copula characterized by the aforementioned marginal distributions and pairwise correlation structure, using the “add_integration” function from the multinma R package.phillippo2024multinma
@@
Finally, our DR augmented MAIC estimator proposed in Section 3.5 produces the estimate: ${\textrm{ATC}}_{DR}=1.332$. Using the non-parametric bootstrap with $B=10000$ resamples, we obtain $\textrm{SE}\left(g(\hat{\mu}_{0}^{1})\right)=0.179$. Then, as per Equation 26, we have $\textrm{SE}\left(\widehat{\textrm{ATC}}_{DR}\right)=0.214$ and a Wald-type 95% CI of $(0.912,1.751)$.
Figure 2 shows the point estimates obtained using the five different estimators alongside their 95% CIs. When comparing the covariate-adjusted approaches to the naïve approach, we observe that covariate adjustment shifts the point estimate towards the null considerably. When comparing normalized IOW, MAIC, G-computation and our proposed DR augmented MAIC estimator, results across the four methods seem consistent. Despite the shift towards the null, results suggest that the intervention improves objective response versus the control, statistically significantly at the 5% level. In this case, the DR point estimate is not meaningfully different than the MAIC or G-computation point estimates, and the DR approach offers slightly increased standard errors and wider CIs than G-computation. Nevertheless, this loss of precision seems a relatively minor price to pay, compensated for by greater reassurance in our results due to increased protection against misspecification of the outcome model.
Figure 2: Point estimates with 95% CIs of the ATC (marginal log-odds ratio of objective response) for the different estimators in the applied example. DR denotes doubly robust and EB denotes entropy balancing.
@@ 6 Discussion
The objective of this article was to clarify existing approaches for doubly robust estimation in the context of externally controlled SATs, and to propose an estimator that augments approaches based on MAIC or entropy balancing via an outcome model. We described and illustrated an extension of this estimator specifically tailored to unanchored ITCs, for the setting with unavailable external control IPD, which is commonly encountered in practice. In a simulation study and applied example, we evaluated the performance and demonstrated the use of different doubly robust augmented estimators, highlighting their merits with respect to the more popular non-augmented singly robust estimators.
@@
Our findings reinforce the understanding that “balancing” approaches to weighting, such as MAIC or entropy balancing, can enhance performance relative to standard “modeling” approaches, but lack the doubly robust property for non-linear outcome models. Conversely, augmented MAIC demonstrates double robustness and exhibits higher precision than non-augmented weighting estimators when the outcome model is correctly specified. Moreover, it generally achieves near-identical precision to G-computation, which offers the lowest variance under correct specification of the outcome model, but may exhibit notable bias where the outcome model is misspecified.
@@
A potential concern about augmented estimators has been possible bias and variance amplification where both the propensity score and the outcome model are incorrectly specified. Nevertheless, such amplification under dual model misspecification was not observed in our simulation study, and one can argue that risks are mitigated by employing “balancing” instead of “modeling” weights (for reasons outlined in the introduction to Section 3.3), but further theoretical work and simulation studies are required to fully support this claim. The results of our simulation study motivate the routine application of doubly robust augmented estimators, particularly those based on MAIC or entropy balancing weights, in practical settings. It is unfortunate to see that virtually all applications of MAIC do not consider augmentation via an outcome model, and that most practical uses of augmented estimators apply “modeling” instead of “balancing” weights.webster2021using
@@
The weighted G-computation estimator described in Section 3.6 exhibited double robustness for the ATC when a canonical logistic link function was defined for the outcome model, but not when the link function was non-canonical, which coincides with what has been suggested previously.gabriel2024inverse Future simulation studies should consider non-binary outcomes, including survival outcomes with censoring, and other summary effect measures beyond the log-odds ratio.
@@
Our simulation study and applied example considered scenarios with relatively low sample sizes in the SAT and the external control, corresponding to typical settings in rare disease and late-stage hematological or solid tumor oncology, where the number of subjects enrolled in SATs can be one- or two-hundred, but may also consist of several hundreds. The number of external controls can be equally small; as such, our findings are potentially applicable where the target estimand is the ATT instead of the ATC, which requires weighting the external control as opposed to the SAT. Somewhat worryingly, all augmented and non-augmented weighting estimators displayed some small-sample bias in our simulation study under a total sample size of $n=200$, even if modeling assumptions were correct. This is consistent with previous research that has also noted small-sample bias.zetterqvist2019doubly, robins2020double
@@
The level of (deterministic) overlap between the SAT and external control covariate distributions in our simulation study was relatively strong. The performance of augmented and non-augmented weighting estimators with respect to G-computation will likely worsen as overlap decreases, particularly in conjunction with small sample sizes. Nevertheless, we conjecture that the performance of balancing-based approaches will suffer less than that of their corresponding modeling-based counterparts, due to generating more stable and less extreme weights. This is unless a complete lack of overlap results in the absence of a solution to the convex optimization balancing problem, in which case extrapolating via G-computation might be the only option.
@@
Finally, it is important to emphasize that all of the covariate-adjusted estimators we considered, including the doubly robust augmented approaches, require the important and untestable assumption of no unmeasured prognostic factors. This is typically the main credibility concern of externally controlled SATs. In practice, important prognostic factors may be unknown or unavailable in at least one of the SAT or external control data sources. Researchers should always consider expert knowledge (e.g., consult clinicians, epidemiologists or statisticians with relevant expertise) and review the relevant literature to evaluate the plausibility of the no unmeasured prognostic factors assumption. An important area of future research is the development of sensitivity analysis or quantitative bias analysis methods to help explore the sensitivity of results to unmeasured prognostic factors, in the specific context of externally controlled SATs and unanchored ITCs.gupta2025quantitative